OpenStudy (anonymous):

divide 10u^2 - 4u / 2u A. 5u - 2 / 2 B. 5u - 2 C. 10u - 4 / 2 D. 10u - 4

OpenStudy (luffingsails):

Hi marie... does the division sign cover both the 10u^2 and 4u term?

OpenStudy (anonymous):

yes!! what does that mean? is there a difference between the division line being short or long?

OpenStudy (luffingsails):

Yes. The answer for what you wrote down is 10u^2-2. But that isn't the right answer.

OpenStudy (luffingsails):

Instead: \[\frac{ 10u^2-4u }{ 2u } = \frac{ 10u^2 }{ 2u } - \frac{ 4u }{ 2u } = 5u - 2\]

OpenStudy (luffingsails):

So each term in the numerator gets divided by the denominator.

OpenStudy (anonymous):

ooooh, so does the longer line mean that the question needs to be simplified/has more steps?

OpenStudy (luffingsails):

Well, it really means that terms are grouped together and needed to be treated as if they were one. So, you can't do something with only one of the terms.

OpenStudy (anonymous):

right, okay. theres another question, it's like the first question: 12mn - 30m / 6m i'm going to try to figure it out on my own, and once i find an answer is it okay if you can tell me if i'm right or if i did something wrong?

OpenStudy (luffingsails):

Sounds good. Better to write it as: (12mn - 30m) / 6m

OpenStudy (anonymous):

would it be 2n - 5?

OpenStudy (luffingsails):

Bingo. Good job.

OpenStudy (anonymous):

omg yay!! thank you lots :)

OpenStudy (luffingsails):

Sure. Have a good morning.